# Simulating Longitudinal Data from Marginal Structural Models

### v1

@@ 5 Demonstration
For subsequent follow-up visits $k>0$:
[…]
$A_{2}$ $=A_{1}=A_{0}$ $A_{k}\sim\operatorname{Bernoulli}(\operatorname{expit}(10-L_{k}+0.1\,A_{k-1}))\text{ for }k=3,4.$
We choose a logistic model for a discrete-time hazard marginal structural model:
@@
where $\operatorname{cum}(\bar{A}_{4})=\Sigma_{i=1}^{4}A_{i}$. This assumes $\exp(\beta)$ as the causal odds ratio of anemia for an incremental dose. If all women receive iron supplements at all four visits, the causal odds ratio would be $\exp(4\beta)$, representing the true value for the “always treated” efficacy estimand.
We use a t-copula $T_{\rho,\operatorname{df}}$ (with $\rho=-0.5$ and $\operatorname{df}=5$) across all pair-copula construction components to model the dependency between $Y$ and each of $L_{0},L_{1},L_{2},L_{3}$, and $L_{4}$. Compared to a Gaussian copula, the t-copula allows for heavier-tailed dependency, capturing stronger associations in the tails. In our example, t-copulas reflect the likelihood that extreme Hb levels (e.g., severely low) are more strongly associated with extreme outcomes (anemia at delivery). This dependency structure may better reflect the underlying association than a Gaussian copula. Our algorithm offers full flexibility, allowing users to choose any copula family and specify different parameters for each bivariate copula in the pair-copula construction.
We simulated scenarios with varying causal coefficients $\beta=\{-0.5,0.3,0\}$, representing strong, moderate and no effect, respectively and considered different samples size $n=\{500,1000,2000\}$. For each combination of $\beta$ and $n$ we generated 400 datasets. We fitted the model $Y=\operatorname{expit}(\alpha+\gamma\,B+\theta\,C-\beta\,\operatorname{cum}(\bar{A}_{4}))$ using two approaches: with stablized inverse probability of treatment weights, and an unweighted approach. For each fit, we computed 95% confidence intervals for $\beta$ using two methods: (1) the robust sandwich estimator of the variance, with rationale provided by Hernán et al. (2000) that the use of weights induces within-subject correlation, and (2) non-parametric bootstrap, constructing 95% confidence intervals from the 2.5% and 97.5% percentiles from 200 bootstrap samples.
Table 1 presents the bias with corresponding Monte Carlo standard errors. The inverse-probability-weighted estimator shows minimal bias, with Monte Carlo standard errors decreasing as sample size increases.In contrast, the unweighted estimator remains significantly biased across sample sizes, showing no convergence. Table 2 reports coverage rates. Both the sandwich and bootstrap CIs approach 95% coverage, though with minor under-coverage. The unweighted estimator, however, exhibits poor coverage due to significant bias in $\beta$. Table 2 shows the statistical power of the test of treatment effect (null hypothesis: $\beta=0$) for each estimator, with power of the weighted estimator increasing as the causal effect size and sample size grow. In summary, these results confirm that data generated using our simulation algorithm accurately reflect the intended causal effects.
@@
Table 1: Bias and Monte Carlo standard errors (in parentheses, defined as $sd(\hat{\beta})$) for inverse-probability-of-treatment-weighted (‘IPTW’) and unweighted estimators (‘unweighted) across different $\beta$ and $n$ values.
@@
Table 2: Coverage of 95% confidence intervals for IPTW-sandwich, IPTW-bootsrap, and unweighted estimators across scenarios with varying $\beta$ and $n$ values.
@@
Table 3: Statistical power of the test for the null hypothesis $\beta=0$, i.e., no causal effect.
@@ 6 Simulating Survival Outcomes
Figure 3: Example directed acyclic graph showing temporal relationships in a longitudinal survival setting. The sequence of outcomes observed over time is denoted as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\dots,K+1\}$.
@@
As shown in Figure 3, we consider the setup described in Section 4 to remain the same except that the outcome now becomes a sequence of outcomes observed over time as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\dots,K+1\}$ and $Y_{k}=1$ represents failure. An example study that follows this setup would be one evaluating the effect of an anti-diabetic drug on cardiovascular risk. The primary endpoint is time to the first occurrence of a major adverse cardiovascular event (MACE) [^2] A composite endpoint of death from cardiovascular causes, non-fatal myocardial infarction, or non-fatal stroke.. The baseline variables include sex and medical history, and an example time-varying confounder might be HbA1c; the treatment influences HbA1c which in turn predicts intercurrent events—for example, use of additional medication—and MACE. We will refer to this example as we explain the algorithm and assumptions in this section.
Similarly, we describe the algorithm for the case where the time-varying confounder is univariate in the absence of any latent procedures. It is fairly straightforward to extend to multiple confounders and latent variables in a manner analogous to ordinary longitudinal outcomes (see Section 4.4).
Depending on the chosen marginal structural model, there are various ways to model the outcomes $Y_{k}$’s. Examples include:
  • Discrete-time Cox marginal structural model, where we model the binary survival outcome $Y_{k}\in\{0,1\}$ directly:
    $\frac{P(Y_{k+1}=1\,|\,do(\bar{a}_{k}),Z,Y_{k}=0)}{P(Y_{k+1}=1\,|\,do(\bar{0}_{k}),Z,Y_{k}=0)}=f(\bar{a}_{k},Z;\alpha).$ (17)
  • Continuous-time hazard MSM, where we model survival time $T$ and observe the binary survival outcome $Y_{k}=I(T>k)$. Specifically
    • Cox proportional hazard: $\lambda(t\,|\,do(\bar{a}_{k}),Z)=\lambda(t\,|\,do(\bar{0}_{k}),Z=z_{0})\exp\{g(\bar{a}_{t},Z;\beta)\}$,
    • Additive hazard: $\lambda(t\,|\,do(\bar{a}_{k}),Z)=\lambda(t\,|\,do(\bar{0}_{k}),Z=z_{0})+g(\bar{a}_{t},Z;\gamma)$.
In Section 6.2 we illustrate the simulation of time-to-event outcomes from a Cox marginal structural model.
@@ 6.1 Simulation algorithm
The general idea is very similar to the non-survival case and the difference is that in the time-step simulation, the simulation of step $k$ is conditional on survival up to the previous time step.
Unlike the derivation of outcome simulation under the frugal parameterization for the non-survival case discussed in Section 4.2, we now need to condition all quantities involved in the copula on survival up to the previous time interval. For simulating survival outcome at the end of the $k$th interval—that is, $y_{k+1}$—(14) becomes:
$p(y_{k+1}$ $\,|\,z,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
$=\prod_{j=0}^{k}c^{*}_{Y_{k+1},L_{j}\mid\bar{L}_{j-1},Z,\bar{A}_{k},Y_{k}=0}\big{\{}F(y_{k+1}\,|\,\bar{l}_{j-1},z,do(\bar{a}_{k}),Y_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},z,do(\bar{a}_{k}),Y_{k}=0);\Theta_{k,j}\}.$ (18)
It is important to note the distinction between the subscripts $j$ and $k$ in the product. To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, that is the $k$th panel in Figure 3. Each copula in (18) associates the outcome $Y_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,1,\dots,k$.
In the non-survival case discussed in Section 4.2, the quantiles, $F(l_{j}\,|\,\bar{l}_{j-1},z,\bar{a}_{j-1})$, for $j=0,1,\dots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation. However, this convenience does not extend when the quantiles are conditioned on survival at later time points. For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim N(0,1)$. Conditioning on survival after 12 months, however, alters the distribution. Survivors’ baseline health likely represents a healthier subgroup than the initial population, meaning the distribution is no longer standard normal. This means that the required conditional quantiles of $L_{k}$ are not directly available and require additional computation. Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
@@
Proposed simulation method: At time step $k$, we specify the following inputs:
$F(L_{k}\,|\,\bar{L}_{k-1},Y_{k}=0)$ $=w_{k,1}$
$F(A_{k}\,|\,\bar{L}_{k},Y_{k}=0)$ $=w_{k,2}$
$F(Y_{k+1}\,|\,\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0)$ $=w_{k,3},$ (19)
where $w_{k,1},w_{k,2},w_{k,3}\sim U(0,1)$. We also have parameters $\Theta_{k,j}$ for each copula $C_{Y_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$ and from the previous step we assume that—by induction—we are given quantiles for each $L_{j}\mid\overline{L}_{j-1},\overline{A}_{j-1},Y_{k-1}=0$. These inputs include the distributions and quantiles of ‘the past’ (i.e. $L_{k}$ and $A_{k}$) given survival, the quantile of the the outcome, $Y_{k+1}$, given the past, and $k+1$ bivariate copulas that ‘weave’ the dependency between the outcome $Y_{k+1}$ and the history of $L$. Crucially, we do not need to compute the full conditional distribution of $Y_{k+1}L_{k}\mid\bar{L}_{k-1},do(\bar{A}_{k}),Y_{k}=0$, but rather just a quantile corresponding to our simulated value; this is central to the efficiency of our approach. We then compute the following for each $j=0,\ldots,k-1$:
@@
$\qquad\quad F(Y_{k}=0\,|\,\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0);\Theta_{k-1,j}\big{\}}$ (20)
$F(Y_{k+1}\,|\,\bar{L}_{j},do(\bar{A}_{k}),Y_{k}=0)$ $=h^{-1}\big{\{}F(Y_{k+1}\,|\,\bar{L}_{j+1},do(\bar{A}_{k}),Y_{k}=0),$
$\qquad\qquad F(L_{j+1}\,|\,\bar{L}_{j},do(\bar{A}_{k}),Y_{k}=0);\Theta_{k,j+1}\big{\}},$ (21)
before
$Y_{k+1}=F^{-1}\big{\{}F(Y_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)\big{\}}.$ (22)
Equation (20) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$. In (21) we iteratively apply $k+1$ $h$-functions to obtain the marginal quantile of $F(Y_{k+1}\,|\,Y_{k}=0)$. Once this quantile is computed, we simulate $Y_{k}$ from the target marginal structural model via inversion, as in (22). In total we evaluate $42k+$ 4 functions (including the simulation of $L_{k}$ and $A_{k}$) for this step.

### v2

@@ 4.4 Illustrative example
where $\operatorname{cum}(\bar{A}_{4})=\Sigma_{i=1}^{4}A_{i}$. Here, $\exp(\beta)$ represents the causal odds ratio for an incremental dose, and $\exp(4\beta)$ corresponds to the “always treated” efficacy estimand.
We use a t-copula $T_{\rho,\operatorname{df}}$ (with $\rho=-0.5$ and $\operatorname{df}=5$) across all pair-copula construction components to model the dependency between $Y$ and each of $L_{0},L_{1},L_{2},L_{3}$, and $L_{4}$. Compared to a Gaussian copula, the t-copula captures heavier-tailed dependency, capturing stronger associations in the tails reflecting stronger associations in extreme scenarios (e.g., severely low hemoglobin levels and anemia). Our algorithm offers full flexibility, allowing users to choose any copula family and specify different parameters for each bivariate copula in the pair-copula construction.
We simulated scenarios with varying causal coefficients $\beta=\{-0.5,-0.3,0\}$, representing strong, moderate and no effect, respectively and samples sizes $n=\{500,1000,2000\}$. For each combination, we generated 400 datasets. The model $Y=\operatorname{expit}(\alpha+\gamma\,B+\theta\,C-\beta\,\operatorname{cum}(\bar{A}_{4}))$ is fitted using stabilized inverse probability of treatment weights, and an unweighted approach. For each fit, we estimated 95% confidence intervals for $\beta$ using both the robust sandwich estimator and non-parametric bootstrap.
Results are summarized in Table 1, 2 and 3. The inverse-probability-weighted estimator shows minimal bias, with magnitude well below Monte Carlo standard errors. In contrast, the unweighted estimator remains significantly biased across sample sizes, showing no convergence. Both the sandwich and bootstrap CIs of the inverse-probability-weighted estimator approach 95% coverage, though with minor under-coverage. The unweighted estimator, however, exhibits poor coverage due to bias. Statistical power for testing the treatment effect (null hypothesis: $\beta=0$) increases with effect size and sample size for the weighted estimator. In summary, these results confirm that data generated using our simulation algorithm accurately reflect the intended causal effects and time-varying confounding is introduced.
@@
Table 1: Bias and Monte Carlo standard errors (in parentheses, defined as $sd(\hat{\beta})$) for inverse-probability-of-treatment-weighted (‘IPTW’) and unweighted estimators (‘unweighted’) across different $\beta$ and $n$ scenarios.
@@
Table 2: Coverage of 95% confidence intervals for IPTW-sandwich, IPTW-bootstrap, and unweighted estimators across scenarios with varying $\beta$ and $n$ values.
@@
Table 3: Statistical power to reject the null hypothesis $\beta=0$; that is, no causal effect.
@@ 5 Simulating Survival Outcomes
Figure 2: Example directed acyclic graph showing temporal relationships in a longitudinal survival setting. The sequence of outcomes observed over time is denoted as $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\ldots,K+1\}$.
@@
In Figure 2, each plate represents a single interval. For the $k$th interval, where $k=0,\ldots,K$, the observations include $L_{k},A_{k}$ at the start of the interval, and the survival status $Y_{k+1}$ at the end. The length of each interval is set to 1. In addition to the notation in Section 4, let $T$ denote an individual’s failure time. The outcome is represented be a sequence $\{Y_{k}\}_{k\in\mathcal{K}^{+}}$, where $\mathcal{K}^{+}=\{1,\ldots,K+1\}$ and $Y_{k+1}=I(T<k+1)$ indicates failure.
We describe the algorithm for the case where the time-varying confounder is univariate and no latent processes are present. The algorithm can be extended analogously to the case of ordinary longitudinal outcomes. The illustrative example at the end of this section demonstrates the simulation with two time-varying confounders.
Depending on the chosen marginal structural model, there are various ways to model the outcomes. Examples include:
  • Discrete-time marginal structural model, where the binary survival outcome $Y_{k}$ is modeled directly:
    $P(Y_{k+1}=1\,|\,do(\bar{A}_{k})=\bar{a}_{k},Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (9)
    where $f$ can be any real-valued functions bounded between 0 and 1, for example, an expit function for a logistic marginal structural model.
  • Continuous-time hazard MSM, where we introduce an auxiliary continuous variable, $\tilde{Y}_{k+1}=T-k\,|\,T>k$, representing the incremental failure time given survival at the start of the $k$th interval. We model incremental survival time $\tilde{Y}_{k+1}$ and truncated it to obtain the observed binary survival outcome $Y_{k+1}$. Given survival at $k$, the time to failure is $T\,|\,T>k=k+\tilde{Y}_{k+1}$. Examples of marginal models for $\tilde{Y}_{k+1}$ include:
    • Cox proportional hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)\exp\{g(\bar{a}_{k},z;\beta)\}$,
    • Additive hazard: $\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),z,y_{k}=0)=\lambda(\tilde{y}_{k+1}\,|\,do(\bar{0}_{k}),z,y_{k}=0)+g(\bar{a}_{k},z;\gamma)$.
We focus on describing the simulation algorithm for time-to-event outcomes, with a remark on adapting the approach for discrete survival outcomes, which follows a similar but simpler framework.
@@ 5.1 Simulation algorithm
The general approach remains similar to the non-survival case, but with the key difference that at each time step $k$, simulation is now conditioned on survival up to the start of the current time interval. For brevity, we continue to omit conditioning on $Z$.
Unlike the derivation of outcome simulation for the non-survival case in Section 4.2, all quantities involved in the copulas must now be conditioned on survival up to this point. Specifically, when simulating the incremental failure time, $\tilde{y}_{k+1}$, given $Y_{k}=0$, (7) becomes:
$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},y_{k}=0)$
$=\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j}\}.$ (10)
It is important to note the distinction between the subscripts $j$ and $k$ in the product. To reiterate, our goal is to simulate the outcome at the end of the $k$th interval, which corresponds to the $k$th panel in Figure 2. Each copula in (10) associates $\tilde{y}_{k+1}$ with all the prior values of the time-varying confounder $L_{j}$, for $j=0,\ldots,k$.
In the non-survival case discussed in Section 4.2, the quantiles $F(l_{j}\,|\,\bar{l}_{j-1},\bar{a}_{j-1})$, for $j=0,\ldots,k$, are directly available at each simulation at step $j$ and can be simply plugged into the iterative application of inverse $h$-functions without further computation. However, this convenience does not extend when the quantiles are conditioned on survival at later time points. For instance, suppose that at $k=0$, we simulate baseline health, $P(L_{0})\sim\operatorname{N}(0,1)$. Conditioning on survival after 12 months, however, alters the distribution. The baseline health of the survivors will almost certainly not be representative of the initial population, meaning that the distribution is no longer standard normal. This means that the conditional quantiles of $L_{j}$ that we need are not directly available and require additional computation. Fortunately, employing copulas to define the dependence structure offers an efficient method to derive these conditional distributions, eliminating the need for resource-intensive techniques such as Monte Carlo approximation or numerical integration.
@@
Proposed simulation method: At time step $k$, we specify the cumulative distribution functions:
$F_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}$ $F_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}$ $F_{\tilde{Y}_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}.$
Additionally, we define $k+1$ bivariate copulas $C_{\tilde{Y}_{k+1}L_{j}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k}=0}$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$. These copulas captures the dependence between $\tilde{Y}_{k+1}$ and each preceding time-varying covariate. Draw $w_{k,1}$, $w_{k,2}$, $w_{k,3}\sim\operatorname{U}(0,1)$ and compute:
$l_{k}=F^{-1}_{L_{k}\mid\bar{L}_{k-1},\bar{A}_{k-1},Y_{k}=0}(w_{k,1}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ $a_{k}=F^{-1}_{A_{k}\mid\bar{L}_{k},\bar{A}_{k-1},Y_{k}=0}(w_{k,2}\,|\,\bar{l}_{k},\bar{a}_{k-1}),$ (11)
to obtain a sample $(l_{k},a_{k})$. We initialize by setting $F(\tilde{y}_{k+1}\,|\,\bar{l}_{k},do(\bar{a}_{k}),y_{k}=0)=w_{k,3}$. Then, for $j=k-1,\ldots,0$, we iteratively apply the transformations:
@@
$\qquad\quad F(\tilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0);\Theta_{{k-1},j}\big{\}}$ (12)
$F(\tilde{y}_{k+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j+1},do(\bar{a}_{k}),y_{k}=0),$
$\qquad\qquad F(l_{j+1}\,|\,\bar{l}_{j},do(\bar{a}_{k}),y_{k}=0);\Theta_{k,j+1}\big{\}}.$ (13)
By induction, the quantiles in the $h$-function in (12) are available from the previous time step $k-1$. We prove this claim as Proposition 5.1.
When the recursion reaches $j=0$, we obtain the quantile $F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0)$. Applying the $h$-function one final time, we set:
$u_{k+1}=F(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}))$ $=h^{-1}\big{\{}F(\tilde{y}_{k+1}\,|\,l_{0},do(\bar{a}_{k}),y_{k}=0),$
$\qquad\qquad F(l_{0}\,|\,do(\bar{a}_{k}),y_{k}=0);\Theta_{k,0}\big{\}}.$
A sample of $\tilde{y}_{k+1}$ is then obtained via inversion:
$\tilde{y}_{k+1}=F_{Y_{k+1}\mid do(\bar{A}_{k}),Y_{k}=0}^{-1}(u_{k+1}),$ (14)
with the corresponding binary outcome $y_{k+1}=I(\tilde{y}_{k+1}<1)$. This completes the $k$th time step, yielding the sample $(l_{k},a_{k},\tilde{y}_{k+1},y_{k+1})$. If $y_{k+1}=1$, the simulation stops; otherwise, proceed to $k+1$.
Crucially, our method avoids computing the full conditional distribution of $\tilde{Y}_{k+1}\mid\bar{L}_{k},do(\bar{A}_{k}),Y_{k}=0$, but rather requiring just a quantile corresponding to our simulated value; this is central to the efficiency of our approach. Equation (12) involves $k$ calculations of the distributions of each $L_{j}\,|\,\bar{L}_{j-1}$ among those who survived to time $k$, in which all the quantities, including quantiles and copulas, are available from the previous step $k-1$. In (13), we recursively apply $k+1$ inverse $h$-functions to obtain the marginal quantile of $F(\tilde{Y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0)$. Finally, we simulate $\tilde{Y}_{k+1}$ from the target marginal structural model via inversion. In total, this time step requires $42k+$ 4 functions, including the simulation of $L_{k}$ and $A_{k}$.